# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ 2.2. Learning and regret
When a buy or sell MO comes, let $(Y_{n})_{n=1}^{N}\in\{0,1\}$ denote the number of the market maker’s filled order and $(\delta_{n})_{n=1}^{N}$ be the offset (depth) posted by the market maker. We know that $\mathcal{L}(Y_{n}|\delta_{n})=\bm{B}(1,e^{-\kappa^{\star}\delta_{n}})$, i.e. the probability of an incoming MO hitting the LO follows the Bernoulli distribution with $p=e^{-\kappa^{\star}\delta_{n}}$.
To learn $\kappa^{\star}$ from the Bernoulli signals, we can simply consider a maximum likelihood estimator [16, Example 7.2.7]. The log-likelihood of $\kappa$ given $(Y_{n})_{n=1}^{N}$ and $(\delta_{n})_{n=1}^{N}$ is
$\ell_{N}(\kappa)=\sum_{n=1}^{N}\left(-\kappa\delta_{n}Y_{n}+(1-Y_{n})\log(1-e^{-\kappa\delta_{n}})\right)\,,$
then
@@
Observe that the value of the solution to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\ell_{N}(\kappa)=0$ may tend to infinity when the number of Bernoulli signals is limited. Therefore, one of the main issues of the likelihood function is that each term in the summation of the Fisher information can be arbitrarily small since $e^{-\kappa\delta}\rightarrow 0$ as $\kappa\rightarrow\infty$ for any fixed $\delta>0$. To overcome this issue, we consider a regularised and biased estimator for $\kappa$. Let us define ~$\tilde{\ell}_{N}(\kappa)$ as a regularised log-likelihood function for estimating $\kappa$ such that
@@
(25) $\begin{split}&\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa)=\bigg{[}\sum_{n=1}^{N}\Big{(}-\delta_{n}Y_{n}+(1-Y_{n})\delta_{n}\frac{e^{-\kappa\delta_{n}}}{1-e^{-\kappa\delta_{n}}}\Big{)}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}\Big{(}-1+\frac{e^{-\kappa\delta_{0}}}{1-e^{-\kappa\delta_{0}}}\Big{)}\bigg{]}\bm{1}_{\kappa\leq\bar{K}}\\ &+\bigg{[}\sum_{n=1}^{N}\Big{(}-\delta_{n}Y_{n}+(1-Y_{n})\delta_{n}\frac{e^{-\bar{K}\delta_{n}}}{1-e^{-\bar{K}\delta_{n}}}\Big{)}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}\Big{(}-1+\frac{e^{-\bar{K}\delta_{0}}}{1-e^{-\bar{K}\delta_{0}}}\Big{)}\bigg{]}\bm{1}_{\kappa>\bar{K}}\\ &+\bigg{[}\sum_{n=1}^{N}\Big{(}-(1-Y_{n})\delta_{n}^{2}\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\Big{)}\bm{1}_{\delta_{n}<+\infty}-\delta_{0}^{2}\frac{e^{-\bar{K}\delta_{0}}}{(1-e^{-\bar{K}\delta_{0}})^{2}}\bigg{]}(\kappa-\bar{K})\bm{1}_{\kappa>\bar{K}},\end{split}$
with $\delta_{0}>0$ a regularisation parameter. In this case, we have
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)=-$ $\left(\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\frac{e^{-\kappa\delta_{n}}}{(1-e^{-\kappa\delta_{n}})^{2}}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}^{2}\frac{e^{-\kappa\delta_{0}}}{(1-e^{-\kappa\delta_{0}})^{2}}\right)\bm{1}_{\kappa\leq\bar{K}}$
$-\left(\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}^{2}\frac{e^{-\bar{K}\delta_{0}}}{(1-e^{-\bar{K}\delta_{0}})^{2}}\right)\bm{1}_{\kappa>\bar{K}}.$
The learning algorithm is summarised as follows.
@@
Choose $\delta_{0}>0$, an initial guess $\kappa_{0}\in[\underline{K},\bar{K}]$, the total number of coming MOs $N$ up to time $T$ with the coming times of the MOs $(t_{n})_{n=1}^{N}$ and the signals of filled LOs from the market maker $(Y_{n})_{n=1}^{N}$, the market maker’s inventory $(Q_{t})_{t\in[0,T]}$
@@
     Choose the offset $\delta_{1}=\psi^{\kappa_{0}}(Q_{0})$ using (18).
@@
     if $\kappa_{i}\notin[\underline{K},\bar{K}]$ then
         Let $\kappa_{i}=\kappa_{i-1}$.
     end if
     Update $\delta_{i+1}=\psi^{\kappa_{i}}(Q_{t_{i}})$ using (18).
@@
We can observe that, compared with a general RL algorithm, there is no trade-off between the exploration and exploitation in Algorithm 1. The agent can always learn from the Bernoulli signals about $\kappa$ since they quote on at least one of the buy side or the sell side.
@@
Step 1: Analysis of Performance Gap
@@
We start with the ergodic analysis for the market making model with misspecified $\kappa$ due to the existence of the term, $\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}},\psi^{\kappa_{t}};\kappa^{\star})\,\mathrm{d}t\Big{]}$, in regret.
@@
(27) $H(q,\delta^{\pm},\bm{p};\kappa^{\star})=\lambda^{+}e^{-\kappa^{\star}\delta^{+}}(p_{1}+\delta^{+})\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\star}\delta^{-}}(p_{2}+\delta^{-})\bm{1}_{q<\bar{q}}-\phi q^{2}\,,$
and the expected reward in the discounted finite-time-horizon setting under model misspecification, $v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\star})$, as
(28) $v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\star})=\mathbb{E}_{q}\bigg{[}\int_{t}^{T}e^{-r(u-t)}f(Q_{u}^{\psi^{\kappa};\kappa^{\star}},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}u+e^{-r(T-t)}G(Q_{T}^{\psi^{\kappa};\kappa^{\star}})\bigg{]}\,,$
where $f$ is given by (21) and $G$ is the terminal condition (11). Then $v_{r}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\star})$ satisfies the following linear ODE. See proof in Appendix A.12.
@@
The function $v_{r}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\star})$ given by (28) satisfies the linear ODE
(29) $0=\partial_{t}v_{r}^{\psi^{\kappa}}-rv_{r}^{\psi^{\kappa}}+H\Big{(}q,\psi^{\kappa},(v_{r}^{\psi^{\kappa}}(t,q^{\prime};T;\kappa^{\star})-v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\star}))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\star}\Big{)}\,,$
@@
Next we focus on the long-run average reward for $v_{0}^{\psi^{\kappa}}$ given by (28) with $r=0$. Proposition 21 provides the existence of $\gamma(\kappa;\kappa^{\star})$, i.e. the long-run average reward for the model with misspecified $\kappa$. Moreover, $\gamma(\kappa;\kappa^{\star})$ with the ergodic value function under the model misspecification, $\hat{v}^{\psi^{\kappa}}(\cdot;\kappa^{\star}):\Omega^{Q}\rightarrow\mathbb{R}$, solves the linear system (31) below. Rigorous definition of $\hat{v}^{\psi^{\kappa}}(\cdot;\kappa^{\star})$ and proof of Proposition 21 are provided in Appendix A.13.
@@
There exists $\gamma(\kappa;\kappa^{\star})\in\mathbb{R}$ such that
(30) $\gamma(\kappa;\kappa^{\star})=\lim_{T\rightarrow+\infty}\frac{1}{T}v_{0}^{\psi^{\kappa}}(0,q;T;\kappa^{\star})\,,$
where $v_{0}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\star})$ is given by (28) with $r=0$. Moreover, there exist $\gamma(\kappa;\kappa^{\star})$ and $\hat{v}^{\psi^{\kappa}}:\Omega^{Q}\rightarrow\mathbb{R}$ that solve the linear system
(31) $0=-\gamma(\kappa;\kappa^{\star})+H\Big{(}q,\psi^{\kappa},(\hat{v}^{\psi^{\kappa}}(q^{\prime};\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star}))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\star}\Big{)},\,\forall q\in\Omega^{Q}.$
The fact that $v_{0}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\star})$ satisfies the linear ODE (29) with $r=0,\,\forall q\in\Omega^{Q}$ allows us to solve it in a matrix form. Moreover, by analysing the case of $T\rightarrow+\infty$ in (30), we can obtain a closed-form expression for $\gamma(\kappa,\kappa^{\star})$ as shown in Proposition 22. See proof in Appendix A.14.
@@
$\bm{\tilde{A}}_{0}(i,q)=\begin{cases}-(\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)}\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)}\bm{1}_{q<\bar{q}}),&\text{if $i=q$},\\ \lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)},&\text{if $i=q+1$},\\ \lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)},&\text{if $i=q-1$},\\ 0,&\text{otherwise, }\end{cases}$
@@
$b_{i}=\lambda^{+}\psi^{\kappa,+}(i)e^{-\kappa^{\star}\psi^{\kappa,+}(i)}\bm{1}_{i>\underline{q}}+\lambda^{-}\psi^{\kappa,-}(i)e^{-\kappa^{\star}\psi^{\kappa,-}(i)}\bm{1}_{i<\bar{q}}-\phi i^{2},$
for $i=[\bar{q},\bar{q}-1,...,\underline{q}]$. Let $\bm{W}$ is the $(\bar{q}-\underline{q}+1)-$square matrix with the first diagonal element equal to $1$ and all other elements equal to 00. Then $\gamma(\kappa;\kappa^{\star})$ given by (30) satisfies
(32) $\gamma(\kappa;\kappa^{\star})\mathds{1}=\bm{U}\bm{W}\bm{U}^{-1}\bm{\tilde{b}},$
@@
Although Proposition 22 gives an expression for $\gamma(\kappa;\kappa^{\star})$, it is not trivial to prove the regularity of $\gamma(\kappa;\kappa^{\star})$ as $\bm{\tilde{A}}_{0}$ is not a self-adjoint or normal operator. We first start with the following lemma that establishes the regularity of $\gamma=\gamma(\kappa)$ given in Theorem 6. The proof is provided in Appendix A.6.
@@
The ergodic constant $\gamma:[\underline{K},\bar{K}]\ni\kappa\mapsto\gamma(\kappa)\in\mathbb{R}$ given by (4) is in $C^{2}([\underline{K},\bar{K}])$.
Next we analyse the regularity of $\gamma(\kappa;\kappa^{\star})$.
@@
$\kappa\mapsto\gamma(\kappa;\kappa^{\star})$ is in $C^{2}([\underline{K},\bar{K}])$.
The key approach in the proof of Lemma 24 (see Appendix A.15) is to construct a self-adjoint operator similar to $\bm{\tilde{A}}_{0}$ and express $\gamma(\kappa;\kappa^{\star})$ in terms of the eigenvector of this self-adjoint operator, which is differentiable.
By Lemma 24, it is trivial to obtain the following lemma by the fact that $\kappa\mapsto\gamma(\kappa;\kappa^{\star})$ attains the maximum at $\kappa=\kappa^{\star}$, because $\psi^{\kappa^{\star}}$ is the optimal control for (30).

### v2

@@ 2.2. Learning and regret
Whenever a MO arrives, the instantaneous fill probability of the market maker’s limit order depends only on the depth (offset) relative to the midprice. The further the market maker’s posted order is from the midprice, the less likely it is to be filled. When a buy or sell MO arrives, let $(Y_{n})_{n=1}^{N}\in\{0,1\}$ denote whether the market maker’s order, posted at depth $(\delta_{n})_{n=1}^{N}$, is filled ($Y_{n}=1$) or not ($Y_{n}=0$). The conditional distribution of $Y_{n}$ given $\delta_{n}$ is modelled as $\mathcal{L}(Y_{n}|\delta_{n})=\bm{B}(1,e^{-\kappa^{\ast}\delta_{n}})$, where $\bm{B}(1,p)$ denotes the Bernoulli distribution and $p=e^{-\kappa\delta_{n}}$ represents the instantaneous fill probability of the market maker’s limit order given a MO arrives.
To learn $\kappa^{\ast}$ from the Bernoulli signals in an online manner, we can simply consider a maximum likelihood estimator [17, Example 7.2.7]. The log-likelihood of $\kappa$ given $(Y_{n})_{n=1}^{N}$ and $(\delta_{n})_{n=1}^{N}$ is
$\ell_{N}(\kappa)=\sum_{n=1}^{N}\left(-\kappa\delta_{n}Y_{n}+(1-Y_{n})\log(1-e^{-\kappa\delta_{n}})\right)\,.$
Clearly
(26) $\frac{\mathrm{d}}{\mathrm{d}\kappa}\ell_{N}(\kappa)=\sum_{n=1}^{N}\left(-\delta_{n}Y_{n}+(1-Y_{n})\delta_{n}\frac{e^{-\kappa\delta_{n}}}{1-e^{-\kappa\delta_{n}}}\right)\,,$
and
@@
However, one may observe that solutions to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\ell_{N}(\kappa_{N})=0$, given by (26), do not necessarily exist. Indeed e.g. if all $Y_{n}=1$ for $n=1$ to $N$ then there is no solution. Moreover, even when a solution $\kappa_{N}$ exists, it may be arbitrarily large making the second derivative $\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\ell_{N}(\kappa_{N})$ arbitrarily small. This is undesirable when quantifying the tail behaviours of the estimator, see also the discussion in Remark 15. To address these issues, we define the regularised log-likelihood function for estimating $\kappa$ by
@@
(27) ~$\begin{split}&\tilde{\ell}_{N}(\kappa)=\big{(}\ell_{N}(\kappa)+R(\kappa)\big{)}\mathds{1}_{\kappa\leq\bar{K}}+\bigg{(}\ell_{N}(\bar{K})+R(\bar{K})\\ &\quad+(\kappa-\bar{K})\Big{(}\tfrac{\mathrm{d}}{\mathrm{d}\kappa}\ell_{N}(\bar{K})+\tfrac{\mathrm{d}}{\mathrm{d}\kappa}R(\bar{K})\Big{)}+\tfrac{1}{2}(\kappa-\bar{K})^{2}\Big{(}\tfrac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\ell_{N}(\bar{K})+\tfrac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}R(\bar{K})\Big{)}\bigg{)}\mathds{1}_{\kappa>\bar{K}}\,.\end{split}$
Recall the assumption that $\bar{K}>\kappa^{\ast}$. The regularisation term $R(\kappa)$ is defined as
$R(\kappa)=-\kappa\delta_{0}+\log(1-e^{-\kappa\delta_{0}})\,,$
where $\delta_{0}>0$ is the regularisation parameter. Observe that as $\bar{K}\rightarrow+\infty$, the regularised log-likelihood ~$\tilde{\ell}_{N}$ converges to $\ell_{N}+R(\kappa)$, i.e. the standard log-likelihood function plus strictly concave regularisation term for any $\delta_{0}>0$. By (27), we have
@@
(28) $\begin{split}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa)&=\left(\frac{\mathrm{d}}{\mathrm{d}\kappa}\ell_{N}(\kappa)+\frac{\mathrm{d}}{\mathrm{d}\kappa}R(\kappa)\right)\mathds{1}_{\kappa\leq\bar{K}}\\ &\hskip 5.0pt+\left(\Big{(}\frac{\mathrm{d}}{\mathrm{d}\kappa}\ell_{N}(\bar{K})+\frac{\mathrm{d}}{\mathrm{d}\kappa}R(\bar{K})\Big{)}+(\kappa-\bar{K})\Big{(}\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\ell_{N}(\bar{K})+\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}R(\bar{K})\Big{)}\right)\mathds{1}_{\kappa>\bar{K}}\,,\end{split}$
and
(29) $\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)=\left(\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\ell_{N}(\kappa)+\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}R(\kappa)\right)\mathds{1}_{\kappa\leq\bar{K}}+\left(\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\ell_{N}(\bar{K})+\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}R(\bar{K})\right)\mathds{1}_{\kappa>\bar{K}}\,.$
Remark 15.
  1. (1)
    By considering the regularised likelihood function ~$\tilde{\ell}_{N}(\kappa)$ (27), we can show that the equation $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa)=0$ always admits a unique solution $\kappa_{N}>0$ for all $N\in\mathbb{N}_{+}$, as stated in Proposition 33. Moreover, we show that any solution $\kappa_{N}>0$ to this equation has the property that $-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa_{N})$ is bounded from below, as stated in Proposition 32. The standard maximum-likelihood estimator does not possess these properties.
  2. (2)
    Note that the depth $\delta_{n}$ posted by the market maker from the ergodic optimal control (19) can take a value of +$\infty$+ ∞ when the inventory hits the boundary. In such cases, the market maker’s order is filled with probability 00, i.e. $Y_{n}=0$ a.s. For the log-likelihood function, we adopt the convention that $0\cdot\infty=0$.
  3. (3)
    Although the regularised estimator guarantees existence, uniqueness and a well-behaved second derivative, the solution to the equation $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=0$ can still be extreme for some $N$. In such cases, the agent’s posted depth $\delta$, determined by the ergodic optimal control (19) as a function of the current inventory, may take values outside of a predefined set $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$ for some constants $\underline{\delta},\bar{\delta}\in\mathbb{R}^{+}$. This boundedness is crucial for establishing the concentration inequality. Furthermore, the second derivative of $\kappa\mapsto\gamma(\kappa;\kappa)$ is not uniformly bounded when $\kappa$ becomes arbitrarily small or large. This property, see Lemma 24, is essential for the second–order performance gap, which leads to a logarithmic regret. Therefore, we impose a constraint on $\kappa_{N}$ in Algorithm 1 to ensure it remains within a compact set. Corollary 35.2 then implies that, with high probability, $\kappa_{N}$ eventually stays within the compact set for all sufficiently large $N$.
The learning algorithm is presented in Algorithm 1.
@@
Choose a small regularisation parameter $\delta_{0}>0$, an initial guess $\kappa_{0}>0$, a truncation function $\varrho(\kappa)=\kappa\mathds{1}_{[\underline{K},\bar{K}]}(\kappa)+\underline{K}\mathds{1}_{(0,\underline{K}]}(\kappa)+\bar{K}\mathds{1}_{[\bar{K},+\infty)}(\kappa)$ with $\underline{K}<\kappa^{\ast}$ and $\bar{K}>\kappa^{\ast}$, the total number $N$ of coming MOs up to time $T$ with the coming times of the MOs $(t_{n})_{n=1}^{N}$ and the signals of filled LOs from the market maker $(Y_{n})_{n=1}^{N}$, the market maker’s inventory $(Q_{t})_{t\in[0,T]}$
@@
     $\hat{\kappa}_{0}=\varrho(\kappa_{0})$
     Choose the offset $\delta_{1}=\psi^{\hat{\kappa}_{0}}(Q_{0})$ using (19).
@@
     $\hat{\kappa}_{i}=\varrho(\kappa_{i})$
     Update $\delta_{i+1}=\psi^{\hat{\kappa}_{i}}(Q_{t_{i}})$ using (19).
@@
In our setting, there is no trade-off between the exploration and exploitation and so Algorithm 1 does not require any exploration phase. This is referred to as the self-exploration property. Even though the agent is exploiting the “optimal” control based on the current estimate of $\kappa$, learning still occurs: whenever a market order arrives, the agent can infer information based on whether its own quote was filled or not, since the agent always quotes on at least one of the buy side or the sell side, i.e. for any $\delta^{\pm}$ take values from the ergodic optimal control (19), we have ($\mathbb{P}\left(\{\delta^{+}=+\infty\}\cap\{\delta^{-}=+\infty\}\right)=0$, ensuring that the agent receives informative feedback over time, which supports the convergence of the estimated parameter to $\kappa^{\ast}$. Of course, for this the assumption that the fill probability parameter $\kappa^{\ast}$ is the same for both buy and sell sides is crucial.
Removing this assumption (that $\kappa^{\ast}$ is the same for both buy and sell sides) would be challenging for two reasons. First, the model would lack explicit solutions. Second, the agent would not be able to keep finite inventory limits while learning.
@@
Step 1: Analysis of the Performance Gap
@@
We start with the ergodic analysis for the market making model with misspecified $\kappa$ due to the existence of the term, $\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}},\psi^{\kappa_{t}};\kappa^{\ast})\,\mathrm{d}t\Big{]}$, in regret.
@@
(31) $H(q,\delta^{\pm},\bm{p};\kappa^{\ast})=\lambda^{+}e^{-\kappa^{\ast}\delta^{+}}(p_{1}+\delta^{+})\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\ast}\delta^{-}}(p_{2}+\delta^{-})\mathds{1}_{q<\bar{q}}-\phi q^{2}\,,$
and the expected reward in the discounted finite-time-horizon setting under model misspecification, $v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\ast})$, as
(32) $v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\ast})=\mathbb{E}_{q}\bigg{[}\int_{t}^{T}e^{-r(u-t)}f(Q_{u}^{\psi^{\kappa};\kappa^{\ast}},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}u+e^{-r(T-t)}G(Q_{T}^{\psi^{\kappa};\kappa^{\ast}})\bigg{]}\,,$
where $f$ is given by (22) and $G$ is the terminal condition (11). Then $v_{r}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\ast})$ satisfies the following linear ODE. See proof in Appendix A.11.
@@
The function $v_{r}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\ast})$ given by (32) satisfies the linear ODE
(33) $0=\partial_{t}v_{r}^{\psi^{\kappa}}-rv_{r}^{\psi^{\kappa}}+H\Big{(}q,\psi^{\kappa},(v_{r}^{\psi^{\kappa}}(t,q^{\prime};T;\kappa^{\ast})-v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\ast}))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\ast}\Big{)}\,,$
@@
Next we focus on the long-term average reward of $v_{0}^{\psi^{\kappa}}$ given by (32) with $r=0$. Proposition 19 provides the existence of $\gamma(\kappa;\kappa^{\ast})$, i.e. the average reward per unit time with a misspecified $\kappa$. Moreover, $\gamma(\kappa;\kappa^{\ast})$ with the ergodic value function under the model misspecification, $\hat{v}^{\psi^{\kappa}}(\cdot;\kappa^{\ast}):\Omega^{Q}\rightarrow\mathbb{R}$, solves the linear system (35) below. Rigorous definition of $\hat{v}^{\psi^{\kappa}}(\cdot;\kappa^{\ast})$ and proof of Proposition 19 are provided in Appendix A.12.
@@
There exists $\gamma(\kappa;\kappa^{\ast})\in\mathbb{R}$ such that
(34) $\gamma(\kappa;\kappa^{\ast})=\lim_{T\rightarrow+\infty}\frac{1}{T}v_{0}^{\psi^{\kappa}}(0,q;T;\kappa^{\ast})\,,$
where $v_{0}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\ast})$ is given by (32) with $r=0$. Moreover, there exist $\gamma(\kappa;\kappa^{\ast})$ and $\hat{v}^{\psi^{\kappa}}:\Omega^{Q}\rightarrow\mathbb{R}$ that solve the linear system
(35) $0=-\gamma(\kappa;\kappa^{\ast})+H\Big{(}q,\psi^{\kappa},(\hat{v}^{\psi^{\kappa}}(q^{\prime};\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast}))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\ast}\Big{)},\,\forall q\in\Omega^{Q}.$
The fact that $v_{0}^{\psi^{\kappa}}(\cdot,\cdot;T;\kappa^{\ast})$ satisfies the linear ODE (33) with $r=0,\,\forall q\in\Omega^{Q}$ allows us to solve it in a matrix form. Moreover, by analysing the case of $T\rightarrow+\infty$ in (34), we can obtain a closed-form expression for $\gamma(\kappa,\kappa^{\ast})$ as shown in Proposition 20. See proof in Appendix A.13.
@@
$\bm{\tilde{A}}_{0}(i,q)=\begin{cases}-(\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}\mathds{1}_{q<\bar{q}}),&\text{if $i=q$},\\ \lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)},&\text{if $i=q+1$},\\ \lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)},&\text{if $i=q-1$},\\ 0,&\text{otherwise, }\end{cases}$
@@
$b_{i}=\lambda^{+}\psi^{\kappa,+}(i)e^{-\kappa^{\ast}\psi^{\kappa,+}(i)}\mathds{1}_{i>\underline{q}}+\lambda^{-}\psi^{\kappa,-}(i)e^{-\kappa^{\ast}\psi^{\kappa,-}(i)}\mathds{1}_{i<\bar{q}}-\phi i^{2},$
for $i=[\bar{q},\bar{q}-1,...,\underline{q}]$. Let $\bm{W}$ is the $(\bar{q}-\underline{q}+1)-$square matrix with the first diagonal element equal to $1$ and all other elements equal to 00. Then $\gamma(\kappa;\kappa^{\ast})$ given by (34) satisfies
(36) $\gamma(\kappa;\kappa^{\ast})\mathds{1}=\bm{U}\bm{W}\bm{U}^{-1}\bm{\tilde{b}},$
@@
Remark 21.
Note that $\gamma(\kappa;\kappa^{\ast})$, given by (36), represents the long-term average reward under the optimal ergodic control with parameter $\kappa$, while the true market environment is $\kappa^{\ast}$. This is different with $\gamma(\kappa)$, which is given by (18). Clearly, $\gamma(\kappa;\kappa)=\gamma(\kappa)$ for any $\kappa\in[\underline{K},\bar{K}]$, meaning that if agent uses the same $\kappa$ as the “true” market parameter, they achieve the optimal long-term average reward. The key challenge is to prove that $[\underline{K},\bar{K}]\ni\kappa\mapsto\gamma(\kappa;\kappa^{\ast})\in\mathbb{R}$ is twice continuously differentiable.
Although Proposition 20 gives an expression for $\gamma(\kappa;\kappa^{\ast})$, it is not trivial to prove the regularity of $\gamma(\kappa;\kappa^{\ast})$ as $\bm{\tilde{A}}_{0}$ is not a self-adjoint or normal operator. We begin with the following lemma that establishes the regularity of $\gamma=\gamma(\kappa)$ given in Theorem 6. This result serves as a preliminary step toward proving Lemma 23. The proof is provided in Appendix A.6.
@@
The ergodic constant $\gamma:[\underline{K},\bar{K}]\ni\kappa\mapsto\gamma(\kappa)\in\mathbb{R}$ given in Theorem 6 is in $C^{2}([\underline{K},\bar{K}])$.
We next analyse the regularity of $\gamma(\kappa;\kappa^{\ast})$.
@@
$\kappa\mapsto\gamma(\kappa;\kappa^{\ast})$ is in $C^{2}([\underline{K},\bar{K}])$.
The key approach in the proof of Lemma 23 (see Appendix A.14) is to construct a self-adjoint operator similar to $\bm{\tilde{A}}_{0}$ and express $\gamma(\kappa;\kappa^{\ast})$ in terms of the eigenvector of this self-adjoint operator, which is differentiable.
By Lemma 23, it is trivial to obtain the following lemma by the fact that $\kappa\mapsto\gamma(\kappa;\kappa^{\ast})$ attains the maximum at $\kappa=\kappa^{\ast}$, because $\psi^{\kappa^{\ast}}$ is the optimal control for (34).